OpenStudy (anonymous):

The perimeter of a rectangle is 62. The area of the same rectangle is 234. Which of the following equations could be used to find the dimensions of the rectangle? A. w(31 – w) = 234 B. w(62 – w) = 234 C. w(31 + w) = 234 D. w(62 + w) = 234

OpenStudy (anonymous):

Can anybody please help?

OpenStudy (anonymous):

Factor out the cheese cake and eat the frosting, then apply Nicholas Cage to both sides

OpenStudy (anonymous):

Simplify kill both the monkeys, and eat the banana

OpenStudy (zpupster):

lw = area and l+l+w+w = 2l+2w = perimeter the answers are in the term of w so 2l + 2w = 62 2l = 62 - 2w l = (62-2w)/2 l = 31 - w substitute that in the equation for area: w(31-w) = 234

OpenStudy (anonymous):

Beat me to it :P

OpenStudy (anonymous):

Thank you so much!!!